You are considering making a movie. The movie is expected to cost $10.7 million up front and take a year to produce. After that, it is expected to make $4.9 million in the year it is released and $1.7 million for the following four years. What is the payback period of this investment? If you require a payback period of two years will you make the movie? Does the movie have positive NPV if the cost of capital is 10.7%

Please go step by step because I'm extremely confused and write legibly. Thank You.
Answer #1

All units in $

Cost in year 0 = 10.7mn

Recovery in year 1 = 4.9 mn

Recovery in year 2 = 1.7 mn

Recovery in year 3 = 1.7 mn

Recovery in year 4 = 1.7 mn

Recovery in year 5 = 1.7 mn

Payback at end of year 4 = 4.9+1.7+1.7+1.7 = 10 mn

Still 10.7-10 = 0.7 mn needs to be recovered which happens in the 5th year.

Time needed = 4 years + (0.7/1.7) of 5th year = 4.41 years

So, if payback is needed in 2 years, one should not take up this project.

NPV = sum of all cashflows discounted

NPV = -10.7 + 1.7 / (1+10.7%)^1 + 1.7 / (1+10.7%)^2+ 1.7 / (1+10.7%)^3+ 1.7 / (1+10.7%)^4 = -5.39 mn

Hence negative NPV